solve y'+y= 1 by power series solution method
1
Expert's answer
2020-06-17T18:58:08-0400

Let power series solution of given differential equation be "y= \sum_{n=0}^{\infin} c_nx^n" .

"\implies y'= \sum_{n=1}^{\infin} nc_nx^{n-1}" .

By putting in given differential equation, we got

"\sum_{n=1}^{\infin}n c_nx^{n-1}+ \sum_{n=0}^{\infin} c_nx^n = 1"

"\implies \sum_{n=0}^{\infin}(n+1) c_{n+1}x^{n}+ \sum_{n=0}^{\infin} c_nx^n = 1 \\
\implies \sum_{n=0}^{\infin}((n+1) c_{n+1}+c_n)x^{n}=1"

Now, by comparing the coefficient on both sides we get,

"c_1+c_0 = 1 , \ and \ (n+1) c_{n+1} + c_n = 0 \ \forall n\geq 2 \implies c_{n+1} = -\frac{c_n}{n+1}"

"\implies c_1=1-c_0, c_2 =- \frac{c_1}{2} = \frac{c_0-1}{2}, c_3 = \frac{c_2}{3} = \frac{c_0-1}{3!}, c_4= \frac{c_0-1}{4!}" and so on.


So, power series solution of given differential equation is

"y= c_0+(1-c_0)x + \frac{c_0-1}{2!} x^2 + \frac{c_0-1}{3!} x^3+ ..."